OpenStudy (anonymous):

The diameter of the circle above is 28 in. What is the circumference of the circle? (Use = 3.14.) A. 43.96 in B. 1,230.88 in C. 87.92 in D. 615.44 in

OpenStudy (kidrah69):

\[C=2\pi~r\]

OpenStudy (kidrah69):

28/2 =radius

OpenStudy (kidrah69):

what is 28/2?

OpenStudy (anonymous):

14

OpenStudy (kidrah69):

now plug that into the formula

OpenStudy (anonymous):

Your answer would be C

OpenStudy (kidrah69):

\[C=2(3.14)~(14)\]

OpenStudy (anonymous):

:) Start with the circumference of a circle formula C=3.14*28 Replace pi with 3.14 (this is approximate) and plug in d=28 (the given diameter)

OpenStudy (anonymous):

ty the algebrahelper but i want to get the explanation

OpenStudy (kidrah69):

i already explained it

OpenStudy (anonymous):

we both did lol

OpenStudy (kidrah69):

C=2(3.14) (14)

OpenStudy (anonymous):

ok mr kidrah69 could you medal thealgebrahelper so then i can medal you n you both get a medal?

OpenStudy (anonymous):

and tyvm

OpenStudy (anonymous):

@kidrah69

OpenStudy (anonymous):

his a lil mad :3 that i gave the answer first .-.

OpenStudy (anonymous):

i gave him a medal

OpenStudy (anonymous):

ok i will medal you n ty

OpenStudy (anonymous):

bye